# Assumption-lean and Data-adaptive Post-Prediction Inference

### v3

@@ Appendix B Proofs
B.1 Proof of Theorem 1
@@
$\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})=\Psi_{n}^{Y}+\bm{\omega}\odot(-\Psi_{n}^{\widehat{f}}+\Psi_{N}^{\widehat{f}})={\bf 0},\text{ where}$ (33)
$\Psi_{n}^{Y}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(y_{i},{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}),\Psi_{n}^{\widehat{f}}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(\widehat{f}({\bf z}_{i}),{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}),\text{ and }\Psi_{N}^{\widehat{f}}=\dfrac{1}{N}\sum_{i=n+1}^{n+N}\psi(\widehat{f}({\bf z}_{i}),{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}).$ (34)
We first establish consistency following Van der Vaart, (2000, Theorem 5.9) by checking its two conditions. The deterministic condition is implied by Condition (C1) and (C2). We are left to verify the uniform convergence condition $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\mathbb{E}(\Psi_{n}^{Y})\|=o_{p}(1)$. By Condition (C3) and Lemma 1, we have
$\|\bm{\omega}\odot\Psi_{n}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|$ $=o_{p}(1),$ (35)
$\|\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}})\|$ $=o_{p}(1),\text{ and}$ (36)
@@
$\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\mathbb{E}(\Psi_{n}^{Y})\|$ (38)
$\leq$ $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})-\Psi_{n}^{Y}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (39)
$=$ $\|-\bm{\omega}\odot\Psi_{n}^{\widehat{f}}+\bm{\omega}\odot\Psi_{N}^{\widehat{f}}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (40)
$\leq$ $\|-\bm{\omega}\cdot\Psi_{n}^{\widehat{f}}+\mathbb{E}(\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|+\|\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}})\|+\|\mathbb{E}(\bm{\omega}\odot\Psi_{N}^{\widehat{f}}-\bm{\omega}\odot\Psi_{n}^{\widehat{f}})\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (41)
@@
which completes the proof of consistency, i.e., $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}_{0}$. Next, we establish the asymptotic normality. Expanding $\Psi(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}})$ at the true value ${\bm{\theta}}_{0}$,
${\bf 0}=$ $\Psi({\bm{\theta}}_{0})+(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}-{\bm{\theta}}_{0})\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}$ (43)
@@
for some $\bar{{\bm{\theta}}}$ between ${\bm{\theta}}_{0}$ and $\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}$. Multiplying both sides with $\sqrt{n}$ and with a direct calculation,
@@
$\left\{\partial\Psi(\bar{\bm{\theta}})/\partial{\bm{\theta}}\right\}^{-1}\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{POP-Inf}}-{\bm{\theta}}_{0})=$ $\sqrt{n}\Psi({\bm{\theta}}_{0})+o_{p}(1).$ (44)
@@
In order to show the asymptotic normality, by Slutsky’s theorem, it’s sufficient to verify the followings:
$\|\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}\|=o_{p}(1),$ (45)
$\sqrt{n}\Psi({\bm{\theta}}_{0})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bf V}(\omega)).$ (46)
For (45),
$\small\|\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}\|\leq$ $\sup_{{\bm{\theta}}\in\Theta}\|\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}\}\|+\|\mathbb{E}\{\partial\Psi_{n}(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}}_{0})/\partial{\bm{\theta}}\}\|.$ (47)
@@
Noticing $\sup_{{\bm{\theta}}\in\Theta}\|\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}})/\partial{\bm{\theta}}\}\|=o_{p}(1)$ by Condition (C4) and Lemma 1 and $\|\mathbb{E}\{\partial\Psi_{n}(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\}-\mathbb{E}\{\partial\Psi_{n}({\bm{\theta}}_{0})/\partial{\bm{\theta}}\}\|=o_{p}(1)$ by $\|\bar{{\bm{\theta}}}-{\bm{\theta}}_{0}\|\leq\|\widehat{\bm{\theta}}-{\bm{\theta}}_{0}\|$ and continuous mapping theorem, equation (45) has been verified. Meanwhile, (46) holds by Central Limit Theorem. The proof of asymptotic normality is completed.
B.2 Proof of Proposition 1

### v4

@@ Appendix B Proofs
B.1 Proof of Theorem 1
We first present a lemma that will be used later.
@@ B.1 Proof of Theorem 1
Condition (C5): There exists a small neighborhood of ${\bm{\theta}}$, denoted as $B$, in which $\bm{\psi}(y,{\bf x},{\bm{\theta}})$ is continuously differentiable with respect to ${\bm{\theta}}$ almost everywhere and the partial derivative satisfies $\mathbb{E}\sup_{{\bm{\theta}}\in B}\lVert\partial\bm{\psi}(Y,{\bf X},{\bm{\theta}})/\partial{\bm{\theta}}\rVert<\infty$ and $\mathbb{E}\sup_{{\bm{\theta}}\in B}\lVert\partial\bm{\psi}(\widehat{f}({\bf Z}),{\bf X},{\bm{\theta}})/\partial{\bm{\theta}}\rVert<\infty$.
@@
$\Psi(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}},{\bm{\omega}})=\Psi_{n}^{Y}+{\rm diag}({\bm{\omega}})(-\Psi_{n}^{\widehat{f}}+\Psi_{N}^{\widehat{f}})={\bf 0},\text{ where}$ (33)
$\Psi_{n}^{Y}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(y_{i},{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}),\Psi_{n}^{\widehat{f}}=\dfrac{1}{n}\sum_{i=1}^{n}\psi(\widehat{f},{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}),\text{ and }\Psi_{N}^{\widehat{f}}=\dfrac{1}{N}\sum_{i=n+1}^{n+N}\psi(\widehat{f},{\bf x}_{i};\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}).$ (34)
We first establish consistency following Van der Vaart, (2000, Theorem 5.9) by checking its two conditions. The deterministic condition is implied by Condition (C1) and (C2). We are left to verify the uniform convergence condition $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}})-\mathbb{E}(\Psi_{n}^{Y})\|=o_{p}(1)$. By Condition (C3) and Lemma 1, we have
$\|{\rm diag}({\bm{\omega}})\Psi_{n}^{\widehat{f}}-\mathbb{E}({\rm diag}({\bm{\omega}})\Psi_{n}^{\widehat{f}})\|$ $=o_{p}(1),$ (35)
$\|{\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}}-\mathbb{E}({\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}})\|$ $=o_{p}(1),\text{ and}$ (36)
@@
$\|\Psi(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}})-\mathbb{E}(\Psi_{n}^{Y})\|$ (38)
$\leq$ $\|\Psi(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}})-\Psi_{n}^{Y}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (39)
$=$ $\|-{\rm diag}({\bm{\omega}})\Psi_{n}^{\widehat{f}}+{\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}}\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (40)
$\leq$ $\|-{\bm{\omega}}\cdot\Psi_{n}^{\widehat{f}}+\mathbb{E}({\rm diag}({\bm{\omega}})\Psi_{n}^{\widehat{f}})\|+\|{\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}}-\mathbb{E}({\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}})\|$ (41)
$+\|\mathbb{E}({\rm diag}({\bm{\omega}})\Psi_{N}^{\widehat{f}}-{\rm diag}({\bm{\omega}})\Psi_{n}^{\widehat{f}})\|+\|\Psi_{n}^{Y}-\mathbb{E}(\Psi_{n}^{Y})\|$ (42)
@@
which completes the proof of consistency, i.e., $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bm{\theta}}$. Next, we establish the asymptotic normality. Expanding $\Psi(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}},{\bm{\omega}})$ at the true value ${\bm{\theta}}$,
${\bf 0}=$ $\Psi({\bm{\theta}},{\bm{\omega}})+(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}-{\bm{\theta}})\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}$ (44)
@@
for some $\bar{{\bm{\theta}}}$ between ${\bm{\theta}}$ and $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}$. Multiplying both sides with $\sqrt{n}$ and with a direct calculation,
@@
$\left\{\partial\Psi(\bar{\bm{\theta}})/\partial{\bm{\theta}}\right\}\sqrt{n}(\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}-{\bm{\theta}})=$ $\sqrt{n}\Psi({\bm{\theta}},{\bm{\omega}}).$ (45)
@@
In order to show the asymptotic normality, by Slutsky’s theorem, it’s sufficient to verify the following:
$\|\partial\Psi(\bar{{\bm{\theta}}},{\bm{\omega}})/\partial{\bm{\theta}}-\mathbb{E}\{\partial\bm{\psi}({\bm{\theta}})/\partial{\bm{\theta}}\}\|=o_{p}(1),$ (46)
$\sqrt{n}\Psi({\bm{\theta}},{\bm{\omega}})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}({\bf 0},{\bf V}({\bm{\omega}})).$ (47)
To show (46), we note that within the neighborhood $B$ in Condition (C5),
$\left\lVert\dfrac{\partial\Psi({\bm{\theta}},{\bm{\omega}})}{\partial{\bm{\theta}}}-\dfrac{\partial\Psi_{n}^{Y}({\bm{\theta}})}{\partial{\bm{\theta}}}\right\rVert$ (48)
$=$ $\left\lVert{\rm diag}({\bm{\omega}})\left(-\partial\Psi_{n}^{\widehat{f}}/\partial{\bm{\theta}}+\mathbb{E}\left[\partial\Psi_{n}^{\widehat{f}}/\partial{\bm{\theta}}\right]\right)+{\rm diag}({\bm{\omega}})\left(\partial\Psi_{N}^{\widehat{f}}/\partial{\bm{\theta}}-\mathbb{E}\left[\partial\Psi_{N}^{\widehat{f}}/\partial{\bm{\theta}}\right]\right)\right\rVert$ (49)
$\leq$ $\left\lVert{\rm diag}({\bm{\omega}})\left(-\partial\Psi_{n}^{\widehat{f}}/\partial{\bm{\theta}}+\mathbb{E}\left[\partial\Psi_{n}^{\widehat{f}}/\partial{\bm{\theta}}\right]\right)\right\rVert+\left\lVert{\rm diag}({\bm{\omega}})\left(\partial\Psi_{N}^{\widehat{f}}/\partial{\bm{\theta}}-\mathbb{E}\left[\partial\Psi_{N}^{\widehat{f}}/\partial{\bm{\theta}}\right]\right)\right\rVert.$ (50)
Then together with Condition (C4) and Lemma 1, we have
$\sup_{\widetilde{\bm{\theta}}\in B}\left\lVert\partial\Psi(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}-\partial\Psi_{n}^{Y}(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}\right\rVert$ $=o_{p}(1).$ (51)
By Conditions (C1, (C4) and Lemma 1, we have
$\sup_{\widetilde{\bm{\theta}}\in B}\left\lVert\partial\Psi_{n}^{Y}(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}-\mathbb{E}\left[\partial\Psi_{n}^{Y}(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}\right]\right\rVert$ $=o_{p}(1).$ (52)
Equations (51) and (52) give $\sup_{\widetilde{\bm{\theta}}\in B}\left\lVert\partial\Psi(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}-\mathbb{E}\left[\partial\Psi_{n}^{Y}(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}\right]\right\rVert=o_{p}(1)$. Then the consistency of $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}$ and the continuous mapping theorem imply that
$\left\lVert\partial\Psi(\bar{{\bm{\theta}}},{\bm{\omega}})/\partial{\bm{\theta}}-\mathbb{E}\left[\partial\bm{\psi}({\bm{\theta}})/\partial{\bm{\theta}}\right]\right\rVert$ (53)
$\leq$ $\sup_{\widetilde{\bm{\theta}}\in B}\left\lVert\partial\Psi(\widetilde{\bm{\theta}})/\partial\widetilde{\bm{\theta}}-\mathbb{E}\left[\partial\Psi_{n}^{Y}({\bm{\theta}})/\partial{\bm{\theta}}\right]\right\rVert+\lVert\mathbb{E}\left[\partial\bm{\psi}(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}\right]-\mathbb{E}\left[\partial\bm{\psi}({\bm{\theta}})/\partial{\bm{\theta}}\right]\rVert$ (54)
$\leq$ $o_{p}(1).$ (55)
@@
Therefore, condition (46) has been verified. Meanwhile, (47) holds by Central Limit Theorem. The proof of asymptotic normality is completed.
B.2 Proof of Corollary 1
We denote the corresponding estimator as $\widehat{\bm{\theta}}(\widehat{\bm{\omega}})$. Then the consistency of $\widehat{\bm{\theta}}(\widehat{\bm{\omega}})$ follows from the consistency of $\widehat{\bm{\omega}}$ and the proof in Section B.1. For asymptotic normality of $\widehat{\bm{\theta}}(\widehat{\bm{\omega}})$, applying Taylor expansion on $0=\Psi(\widehat{\bm{\theta}},\widehat{\bm{\omega}})$ yields
00 0 $=\Psi(\widehat{\bm{\theta}},\widehat{\bm{\omega}})=\Psi({\bm{\theta}},{\bm{\omega}}^{\textnormal{opt}})+{\rm diag}(\widehat{\bm{\omega}}-{\bm{\omega}}^{\textnormal{opt}})\left(-\Psi_{n}^{\widehat{f}}+\Psi_{N}^{\widehat{f}}\right)+(\widehat{\bm{\theta}}-{\bm{\theta}})\partial\Psi(\bar{{\bm{\theta}}})/\partial{\bm{\theta}}$ (56)
for some $\bar{{\bm{\theta}}}$ between ${\bm{\theta}}$ and $\widehat{\bm{\theta}}(\widehat{\bm{\omega}})$. With the same proof technique in Section B.1, we have
$\sqrt{n}(\widehat{\bm{\theta}}(\widehat{\bm{\omega}})-{\bm{\theta}})\stackrel{{\scriptstyle D}}{{\rightarrow}}N({\bf 0},{\bm{\Sigma}}({\bm{\omega}}^{\textnormal{opt}})),$ (57)
which implies Corollary 1.
B.3 Proof of Proposition 1
The $j$-th diagonal element of the asymptotic covariance matrix ${\bm{\Sigma}}({\bm{\omega}})$ of $\widehat{\bm{\theta}}_{\textnormal{{{PSPA}}}}({{\bm{\omega}}})$ given weighting vector ${\bm{\omega}}$ is
${\bm{\Sigma}}_{jj}({\bm{\omega}})=\omega_{j}^{2}[{\bf A}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}^{-1}]_{jj}-2\omega_{j}[{\bf A}^{-1}{\bf M}_{4}{\bf A}^{-1}]_{jj}+[{\bf A}^{-1}{\bf M}_{1}{\bf A}^{-1}]_{jj}.$ (58)
It is a quadratic function of $\omega_{j}$ with a unique minimizer
$\omega_{j}^{\textnormal{opt}}=\dfrac{[{\bf A}^{-1}{\bf M}_{4}{\bf A}^{-1}]_{jj}}{[{\bf A}^{-1}({\bf M}_{2}+\rho{\bf M}_{3}){\bf A}^{-1}]_{jj}}.$ (59)
Therefore, given any weighting vector ${\bm{\omega}}^{*}=[\omega_{1}^{*},\cdots,\omega_{q}^{*}]^{\rm T}$, ${\bm{\Sigma}}_{jj}({\bm{\omega}}^{\textnormal{opt}})\leq{\bm{\Sigma}}_{jj}({\bm{\omega}}^{*})$, which completes the proof.
B.4 Proof of Proposition 2
@@
B.5 Lemma for PSPA when applied to ML-predicted covariates
Lemma 2.
Letting ${\bf A}$ and ${\bf B}$ be two $q\times q$ gram matrices, and ${\bm{\omega}}$ is a q-dimensional vector. If $0\leq\omega_{j}\leq\dfrac{\lambda_{{\min,+}}({\bf A})}{\lambda_{\max}({\bf B})}$ for all $j\in\{1,\cdots,p\}$ then ${\bf A}-\text{diag}({\bm{\omega}}){\bf B}$ is positive semi-definite.
To prove the lemma, we will verify that ${\bf A}-\operatorname{diag}({\bm{\omega}}){\bf B}$ is positive semi-definite by showing that for any vector ${\bf x}$, the quadratic form ${\bf x}^{T}({\bf A}-\operatorname{diag}({\bm{\omega}}){\bf B}){\bf x}\geq 0$.
Let ${\bf x}$ be any vector in $\mathbb{R}^{p}$. Since
${\bf x}^{\rm T}{\bf A}x$ $\geq\lambda_{\min,+}({\bf A})\|{\bf x}\|^{2}$ (73)
${\bf x}^{\rm T}\operatorname{diag}({\bm{\omega}}){\bf B}{\bf x}$ $\leq\lambda_{\max}(\operatorname{diag}({\bm{\omega}}){\bf B})\|x\|^{2}\leq\max_{j}\omega_{j}\lambda_{\max}{\bf B}\|{\bf x}\|^{2}$ (74)
Therefore
${\bf x}^{\rm T}{\bf A}{\bf x}-{\bf x}^{\rm T}\operatorname{diag}({\bm{\omega}}){\bf B}\geq(\lambda_{\min,+}({\bf A})-\max_{j}\omega_{j}\lambda_{\max}{\bf B})\|{\bf x}\|^{2}$ (75)
To ensure ${\bf x}^{\rm T}{\bf A}{\bf x}-{\bf x}^{\rm T}\operatorname{diag}({\bm{\omega}}){\bf B}\geq 0$, we require $\lambda_{\min,+}({\bf A})-\max_{j}\omega_{j}\lambda_{\max}({\bf B})\geq 0$, that is
$\max_{j}\omega_{j}\leq\dfrac{\lambda_{\min,+}({\bf A})}{\lambda_{\max}({\bf B})}$ (76)
B.6 Proof of Corollary 5
The $j$-th diagonal element of the asymptotic covariance matrix ${\bm{\Sigma}}^{\prime}({\bm{\omega}})$ of $\widehat{\bm{\theta}}_{\textnormal{{PSPA}}^{\prime}}$ given weighting vector ${\bm{\omega}}$ is
${\bm{\Sigma}}^{\prime}_{jj}({\bm{\omega}})$ $=\omega_{j}^{2}[{\bf A}^{-1}({\bf M}_{2}^{\prime}+\rho{\bf M}_{3}^{\prime}){\bf A}^{-1}]_{jj}-2\omega_{j}[{\bf A}^{-1}{\bf M}_{4}^{\prime}{\bf A}^{-1}]_{jj}+[{\bf A}^{-1}{\bf M}_{1}^{\prime}{\bf A}^{-1}]_{jj},$ (77)
It is a quadratic function of $\omega_{j}$ with a unique minimizer
$\omega_{j}^{\textnormal{opt},*}=\dfrac{[{\bf A}^{-1}{\bf M}^{\prime}_{4}{\bf A}^{-1}]_{jj}}{[{\bf A}^{-1}({\bf M}^{\prime}_{2}+\rho{\bf M}^{\prime}_{3}){\bf A}^{-1}]_{jj}}.$ (78)
If $\omega_{j}^{\textnormal{opt},*}>0$, ${\bm{\Sigma}}^{\prime}_{jj}({\bm{\omega}})$ is a decreasing function with $\omega_{j}\in[0,\omega_{j}^{\textnormal{opt},*}]$. If $\omega_{j}^{\textnormal{opt},*}\leq 0$, ${\bm{\Sigma}}^{\prime}_{jj}({\bm{\omega}})$ ${\bm{\Sigma}}^{\prime}_{jj}({\bm{\omega}})$ is an increasing function with $\omega_{j}\in[\omega_{j}^{\textnormal{opt},*},0]$. We also have $\frac{\lambda_{{\min,+}}[\frac{1}{n}\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})]}{\lambda_{\max}[\frac{1}{n}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]}>0$
Therefore, given $\hat{\bm{\omega}}^{\textnormal{opt}}=[\widehat{\omega}_{1}^{\textnormal{opt}},\dots,\widehat{\omega}_{q}^{\textnormal{opt}}]^{\rm T}$, where
$\widehat{\omega}_{j}^{\textnormal{opt}}=\begin{cases}\min(\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}},\frac{\lambda_{{\min,+}}[\frac{1}{n}\nabla\bm{\psi}(y_{i},{\bf x}_{i};{\bm{\theta}})]}{\lambda_{\max}[\frac{1}{n}\nabla\bm{\psi}(y_{i},\widehat{q};{\bm{\theta}})]})&\quad\text{if }\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}}>0\\ \max(\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}},0)&\quad\text{if }\frac{[\widehat{\bf A}^{-1}\widehat{\bf M}_{4,{\text{C}}}^{{}^{\prime}}\widehat{\bf A}^{-1}]_{jj}}{[\widehat{\bf A}^{-1}(\widehat{\bf M}_{2,{\text{C}}}^{{}^{\prime}}+\rho\widehat{\bf M}_{3,{\text{C}}}^{{}^{\prime}})\widehat{\bf A}^{-1}]_{jj}}\leq 0\\ \end{cases}$ (79)
for all $j\in 1,\dots,q$, and the weighting vector ${\bm{\omega}}^{\textnormal{C}}=[0,\cdots,0]^{\rm T}$ that correspond to the classical estimator, ${\bm{\Sigma}}_{jj}({\bm{\omega}}^{\textnormal{opt}})\leq{\bm{\Sigma}}_{jj}({\bm{\omega}}^{\textnormal{C}})$, which completes the proof.